The resistant (R) of a given coil of wire per 1km is 0.4ohms. If 40 of the wire is used to supply 16 amps to an 8kv rated x ray generator connected to 420 v mains supply, calculate the voltage drop on load.
1
Expert's answer
2018-09-24T10:41:08-0400
The 8 kV rated X-ray generator, is fed by V_ps=420 V power supply, which requires I=16 A current. There is 40 without measure unit in the condition; assume that it means 40 metres. Calculate the resistance of the wire:
R_w=40/1000·0.4=0.016 Ω.
Now calculate the voltage drop across the load (i.e. generator) by Kirchhoff’s voltage law:
V_ps=IR_g+IR_w,
IR_g=V_ps-IR_w=420-16·0.016=419.74 V.